Ahmed’s Company manufactures and sells one product. The product has the following cost and revenue data:
                                                                                   

Selling price Per Unit (AED) 70
Variable cost Per Unit (AED) 25



Total fixed expenses per month are as follows:

Expenses types AED
Advertising 200,000
Rent 100,000
Heating 100,000



The company produced and sold 10,000 units during the month and had no beginning or ending inventories.
Match the closest correct answers for the below questions:

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.   

What is the break-even point in dirhams?

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.   

How many units would the company have to sell to have a profit of AED120,000?

      -       A.       B.       C.       D.       E.       F.       G.       H.       I.       J.   

If the sales increase by 200 units, by how much should the net operating income (profit) increase?

A.

[ AED 30,000]

B.

[ AED 110,000]

C.

[10,800 UNITS]

D.

[ AED 400,000]

E.

[12,000 UNITS]

F.

[11,556 UNITS]

G.

[40,000 UNITS]

H.

[AED 9,000]

I.

[AED 622,222]

J.

[ AED 50,000]
Answer #1

Total Fixed cost (TFC) = Advertising expense + Rent expense + Heating expense = 200,000 + 100,000 + 100,000 = 400,000

Contribution margin (CM) (AED) = Unit Selling price - Unit variable cost = 70 - 25 = 45

(1) Option (I)

BEP in dirhams = TFC / (CM / Selling price) = 400,000 / [45 / 70) = 2,800,000 / 45 = 622,222

(2) Option (F)

BEP in units = (TFC + Profit) / CM = (400,000 + 120,000) / (70 - 25) = 520,000 / 45 = 11,556

(3) Option (H)

Initial Sales (Q) = 10,000 units.

Initial Profit = (Q x CM) - FC = (10,000 x 45) - 400,000 = 450,000 - 400,000 = 50,000

New sales (Q) = 10,000 + 200 = 10,200

New Profit = (10,200 x 45) - 400,000 = 459,000 - 400,000 = 59,000

Profit has increased by AED (59,000 - 50,000) = 9,000